# Understanding Before Reasoning: Enhancing Chain-of-Thought with Iterative Summarization Pre-Prompting

### v1

@@ 4.3 Ablation Studies
Table 5: Performance comparison of $\text{ISP}^{2}$ aiding CoT reasoning under different combinations and summarization order settings.
@@
Table 6: Performance comparison of $\text{ISP}^{2}$ across different combinations and ordering of summarization steps, where Self Consistency @5 is utilized by retrieving five chains of CoT reasoning to make majority votes.
@@
Table 7: Performance comparison of $\text{ISP}^{2}$ aiding Complex CoT reasoning under different combinations and summarization order settings.
@@
Table 8: Performance comparison of $\text{ISP}^{2}$ across different combinations and ordering of summarization steps, where Self Consistency @5 is utilized by retrieving five chains of Complex CoT reasoning to make majority votes.
Information pair extraction and two low-score information pair summarization are effective for answering questions. In tests across six datasets, both CoT and Complex CoT performance improved with the addition of entities and complete information pairs. However, the inclusion of complete information pairs was more significantly beneficial for problem reasoning. Information pairs offer more detailed explanations compared to entity extraction. While entity extraction helps LLMs focus on key terms, OE reasoning requires the model to balance understanding the context and solving the problem simultaneously, which can complicate the reasoning process. On the other hand, OIP reasoning provide detailed descriptions that simplify the reasoning process by allowing the LLM to concentrate solely on solving the problem, thereby reducing unnecessary complexity and enhancing effectiveness. Additionally, when $\text{ISP}^{2}$ switched to extracting the two highest scoring information pairs for iterative summarization, the results were not competitive. The highest scores indicate that the information pairs already possess a significant positive boost to answering the question. After summarization, further exploration of in-depth information becomes limited. In contrast, low-scoring information pairs, having been retained through the filtering process, prove their efficacy yet still have substantial potential to be tapped. Therefore, choosing two low-scoring information pairs allows for the accumulation of more implicit knowledge through iterative summarization, enabling the information pairs to continuously deepen their cognitive approach towards providing more reliable knowledge.
5 Discussion
@@
5.1 Summarization Steps Analysis
5.2 Error Source Analysis
We extracted 100 erroneous samples from each dataset and identified three critical types of errors that arise during the $\text{ISP}^{2}$ process. 1. Information Pair Error (IPE): Failure to extract all explicit information pairs in the question or extraction of information pairs with misleading information; 2. Summarization Error (SE): Summaries that contain misleading information; 3. Reasoning Error (RE): Correct summarization but generation of an incorrect answer due to faulty reasoning. These types of errors represent potential points of failure at each step of the $\text{ISP}^{2}$ process, which can propagate and affect subsequent operations.
@@ 5.2 Error Source Analysis
Table 9: Proportion of different error categories across various datasets.
Table 9 displays the distribution of error categories across each dataset. Considering the error categories, the probability of information pair extraction errors is the highest, while the error rate for summarization is relatively lower, and the error rate for inference is the lowest. We observe that in commonsense reasoning datasets, the error rates for information extraction and summarization are close. It is attributed to the fact that the model’s own knowledge contains some elements related to implicit information, so the defects in information pair extraction are not significant. However, during the final inference stage, conflicts often arise between the LLM’s inherent knowledge and the already summarized information, leading to judgment errors in the final decision-making process. The rate of information pair extraction errors in mathematical datasets shows a positive correlation with the complexity of the dataset. From basic arithmetic operations to complex algebraic calculations, this trend becomes increasingly pronounced. It indicates that as the difficulty of problems increases, the LLM’s limitations in mathematical understanding become more apparent, leading to a greater impact on its ability to accurately collect and process mathematical information. Additionally, summarization helps improve dataset accuracy, but deficiencies in the large model’s mathematical computation capabilities during the summarization process still exist. Indeed, this issue can also arise during the inference process. Even if the summarized content is accurate, computational errors can still occur during inference.
6 Conclusion
@@
7 Appendix
[…]
Answer Formation
Here is the $\text{ISP}^{2}$ process dialogue, with different steps handled through CoT. Table 11 and Table 12 show the reasoning processes for mathematical and commonsense problems, respectively. In the final part of our reasoning to generate the answer, we will fix the format of the final answer, as shown in Table 10.
Tasks Answer-format Instructions
gsm8k Your final answer should be a single numerical number, in the form answer.
svamp Your final answer should be a single numerical number, in the form answer.
AQuA Your answer should be a single numerical number in the form answer,
and there should be only one correct choice in the form choice.
AddSub Your final answer should be a single numerical number, in the form answer.
CommonsenseQA Your answer should be in the form choice. There is only one correct choice.
StrategyQA Your answer should be either yes or no, in the form answer.
Table 10: Answer format instructions for various tasks
SVAMP Problem
Q: Kylar went to the store to buy glasses for his new apartment. One glass costs \$5, but every second glass costs only 60% of the price. Kylar wants to buy 16 glasses. How much does he need to pay for them?
Adaptive Extraction of Candidate Information
Information Pairs
Glasses
- One glass costs \$5.
- Every second glass costs 60% of the price.
- Kylar wants to buy 16 glasses in total.
Kylar
- Kylar went to the store.
- Wants 16 glasses.
Price
- One glass costs \$5.
- Every second glass costs 60%.
- Kylar needs to pay for all glasses.
Iterative Summarization for Knowledge Understanding
Step1
[…]
- Buying 16 glasses.
- One glass \$5, second glass 60% of that.
Step2
[…]
- 16 glasses for new apartment.
- Alternating prices, \$5 and \$3.
- Calculate total based on pricing scheme.
Final Input Formation
Q: Kylar went to the store to buy glasses for his new apartment. One glass costs \$5, but every second glass costs only 60% of the price. Kylar wants to buy 16 glasses. How much does he need to pay for them?
Glasses Purchase
- Kylar is purchasing 16 glasses for his new apartment.
- The price of one glass is \$5, with every second glass costing only 60% of the original price.
- Kylar needs to pay for all the glasses he wants to buy.
- The total cost of the glasses can be calculated based on the pricing scheme.
A: Let’s think step by step.
Output
Kylar wants to buy 16 glasses for his new apartment.
The price of the first glass is \$5.
The price of every second glass is 60% of the original price, which is $0.6\times 5=3$.
This creates an alternating pattern of $5$ and $3$ for each glass.
So, for every pair of glasses, Kylar pays \$5 + \$3 = \$8.
Since there are 16 glasses in total, there are $\frac{16}{2}=8$ pairs of glasses.
[…]
Answer: 64
Table 11: Inference Process of $\text{ISP}^{2}$ on Mathematical Problem Datasets.
SQA Problem
Q: Do the anchors on Rede Globo speak Chinese?
Adaptive Extraction of Candidate Information
Information Pairs
Anchors on Rede Globo
- The anchors on Rede Globo refer to the news presenters or hosts on the Brazilian television network Rede Globo, one of the largest TV networks in Brazil.
- These anchors are responsible for delivering news, conducting interviews, and hosting various programs on the network.
- Rede Globo anchors are typically well-versed in multiple languages to cater to a diverse audience.
Chinese language
- Chinese is one of the most widely spoken languages globally, with numerous dialects such as Mandarin and Cantonese.
- Learning Chinese can be challenging due to its tonal nature and complex writing system.
- It is unlikely that anchors on Rede Globo speak Chinese unless they have a specific background or training in the language.
Iterative Summarization for Knowledge Understanding
Step1
[…]
- The anchors on Rede Globo, prominent figures on the Brazilian television network, are adept at presenting news, conducting interviews, and hosting programs for a diverse audience.
- Given the global reach of Rede Globo, these anchors are likely proficient in multiple languages to cater to a wide viewership.
- While Chinese is a widely spoken language with complexities in tonality and writing, it is improbable that anchors on Rede Globo speak Chinese unless they possess specialized training or background in the language.
Final Input Formation
Q: Do the anchors on Rede Globo speak Chinese?
Anchors on Rede Globo
- The anchors on Rede Globo, prominent figures on the Brazilian television network, are adept at presenting news, conducting interviews, and hosting programs for a diverse audience.
- Given the global reach of Rede Globo, these anchors are likely proficient in multiple languages to cater to a wide viewership.
- While Chinese is a widely spoken language with complexities in tonality and writing, it is improbable that anchors on Rede Globo speak Chinese unless they possess specialized training or background in the language.
A: Let’s think step by step.
Output
Anchors on Rede Globo are skilled in presenting news for a diverse audience.
They likely know multiple languages to cater to a wide viewership.
Chinese is a complex language with tonality and writing systems.
It is improbable that anchors on Rede Globo speak Chinese unless they have specialized training in the language. So the answer is no.
Answer: No
Table 12: Inference Process of $\text{ISP}^{2}$ on Commonsense Problem Datasets
[…]
Evaluation Metrics
Accuracy is used as an evaluation metric to measure performance in tasks involving mathematical reasoning and commonsense reasoning. For problems where the output is a number, we use regular expressions for exact match evaluation for datasets such as GSM8K, SVAMP, and AddSub. However, in AddSub, we find that LLMs cannot perform computations precisely (especially when dealing with high-precision floats and large numbers), so we adopt the method with a relative tolerance of 0.001 to compare answers. For AQuA, since it involves mathematical multiple-choice questions, we compare the extracted option with the correct option to verify consistency, to prevent incorrect problem-solving occurrences. In cases where the commonsense dataset answers are binary (yes / no) or multiple-choice, such as StrategyQA / CommonsenseQA, we assess whether the extracted result aligns with the provided label.
Acknowledgements.
This work was supported in part by the Science and Technology Commission of Shanghai Municipality under Grant (21DZ2203100), in part by the National Natural Science Foundation of China under Grant (62006150), in part by Shanghai Local Capacity Enhancement project (21010501500) and in part by Science and Technology Innovation Action Plan of Shanghai Science and Technology Commission for social development project under Grant (21DZ1204900).
[…]

### v2

@@ IV-C Ablation Studies
TABLE V: Ablation study on $\text{ISP}^{2}$ aiding CoT reasoning using Self-Consistency.
@@
Information pair extraction and two low-score information pair summarization are effective for answering questions. In our experiments across six datasets, both CoT and Complex CoT methods showed performance improvements when entities and complete information pairs were incorporated. The inclusion of complete information pairs (OIP) provided significantly greater benefits for problem reasoning compared to entity extraction (OE). OIP provides a structured contextual description that enables the model to directly focus on the problem-solving logic without being distracted by additional context understanding. In contrast, OE merely extracts scattered keywords or entities, forcing the model to reconstruct fragmented information during reasoning. OE not only increases cognitive load but also makes the model prone to errors due to misinterpretation of the context.
The most intriguing finding was the underperformance of ISP2 when selecting the two highest-scoring information pairs for iterative summarization. This counter-intuitive result can be attributed to the inherent limitations of high-scoring pairs. While these pairs are prioritized for their surface-level relevance, their utility often saturates early during summarization. For instance, in mathematical reasoning tasks, high-scoring pairs might repeatedly emphasize a single solution path without addressing edge cases or alternative formulations. The early convergence leaves little room for deeper exploration, ultimately restricting the model’s ability to refine its reasoning.
In contrast, selecting low-scoring information pairs yielded superior results. These pairs, though initially deemed less relevant, often contain indirect or non-obvious associations. Their diversity introduces latent constraints or cross-domain connections that enrich the reasoning process. Iterative summarization of low-scoring pairs creates a self-correcting feedback loop: if one pair suggests an erroneous solution path, subsequent pairs may introduce conflicting evidence, prompting the model to reconcile discrepancies. The dynamic mimics human-like critical thinking, where hypotheses are refined through incremental validation. High-scoring pairs reduce uncertainty prematurely, increasing the risk of overcommitting to incorrect premises. Low-scoring pairs, by preserving diverse perspectives, act as a “checks-and-balances” system that balances exploitation of strong signals with exploration of implicit knowledge. The selection mechanism of low-score pair is particularly vital in complex tasks where superficial relevance must be weighed against deeper logical consistency.
TABLE VI: Ablation study on $\text{ISP}^{2}$ aiding Complex CoT reasoning.
@@
TABLE VII: Ablation study on $\text{ISP}^{2}$ aiding Complex CoT reasoning using Self-Consistency.
@@
V Discussion
V-A Summarization Steps Analysis
@@
V-B Error Source Analysis
We extract 100 erroneous samples from each dataset and identified three critical types of errors that arise during the $\text{ISP}^{2}$ process. 1. Information Pair Error (IPE): Failure to extract all explicit information pairs in the question or extraction of information pairs with misleading information; 2. Summarization Error (SE): Summaries that contain misleading information; 3. Reasoning Error (RE): Correct summarization but generation of an incorrect answer due to faulty reasoning. These types of errors represent potential points of failure at each step of the $\text{ISP}^{2}$ process, which can propagate and affect subsequent operations.
TABLE VIII: Proportion of different error categories across various datasets.
@@ V-B Error Source Analysis
Table VIII displays the distribution of error categories across each dataset. Considering the error categories, the probability of information pair extraction errors is the highest, while the error rate for summarization is relatively lower, and the error rate for inference is the lowest. We observe that in commonsense reasoning datasets, the error rates for information extraction and summarization are close. It is attributed to the fact that the model’s own knowledge contains some elements related to implicit information, so the defects in information pair extraction are not significant. However, during the final inference stage, conflicts often arise between the LLM’s inherent knowledge and the already summarized information, leading to judgment errors in the final decision-making process. The rate of information pair extraction errors in mathematical datasets shows a positive correlation with the complexity of the dataset. From basic arithmetic operations to complex algebraic calculations, this trend becomes increasingly pronounced. It indicates that as the difficulty of problems increases, the LLM’s limitations in mathematical understanding become more apparent, leading to a greater impact on its ability to accurately collect and process mathematical information. Additionally, summarization helps improve dataset accuracy, but deficiencies in the large model’s mathematical computation capabilities during the summarization process still exist. Indeed, this issue can also arise during the inference process. Even if the summarized content is accurate, computational errors can still occur during inference.
VI Conclusion
@@
VII Acknowledgments
This work was supported in part by the Science and Technology Commission of Shanghai Municipality under Grant (21DZ2203100), in part by the National Natural Science Foundation of China under Grant (62006150).